Saturday 18 September 1999 9.00am EST
|During the last 24 hours Chandra exited radiation belts and a post belt ACIS calibration measurement was performed. HRC/LETG calibration observations continued until 9AM 9/18. The grating insertion following belt exit was monitoried in real-time and proceeded nominally. A memory dump of the grating parameters verified the correct states had been set.
The ACIS diagnostic load was completed and reviewed over night. The load includes operation of focal plane at -60 deg C with the internal calibration source. This measurement is being made to distinguish between possible effects causing the degraded energy resolution of the front-side illuminated chips. The last internal cal source measurement indicated that the front-side chips are showing continued energy degradation, but that the back-side chips remain unchanged.
The ACIS diagnostic measurements will continue for the next 24 hours leading up to the radiation belt entry. Following a pre-belt calibration measurement with ACIS, the Observatory will be placed in its eclipse attitude and pass through a 1 hour eclipse (penumbra and umbra) starting at 2PM 9/19. During the next belt passage the ACIS will be placed at the out-of focus position.
Return to Status Reports